Ziphip North Burials Groww x C Situtional Software Ca x Mind Tup-Cengage Leaming x Farel Online Student Work X + Club Peety SVideo Excel Online Structured Activity: Hamada equation Situational Software Co. (SSC) is trying to establish its optimal capSituational Software Co. (SSC) is trying to establish its optimal capital structure. Its current capital structure consists of 35% debt and 65% equity; however, the CEO believes that the firm should use more debt. The risk-free rate, rRF, is 6%; the market risk premium, RPM, is 7%; and the firm's tax rate is 40%. Currently, SSC's cost of equity is 13%, which is determined by the CAPM. The data has been collected in the Microsoft Excel Online file below. Open the spreadsheet and perform the required analysis to answer the question below.

What would be SSC's estimated cost of equity if it changed its capital structure to 50% debt and 50% equity? Round your answer to two decimal places. Do not round intermediate steps.
Answer #1

First, we need to find the current beta;

Cost of equity = Risk-free rate + [Beta * Market Risk Premium]

13% = 6% + [Beta * 7%]

13% - 6% = Beta * 7%

Beta = 7% / 7% = 1

Now, we will find the unlevered beta

Unlevered Beta = Levered Beta / [1 + {(1 - t) * (D/E)}]

= 1 / [1 + {(1 - 0.40) * (0.35 / 0.65)}]

= 1 / [1 + 0.3231] = 1 / 1.3231 = 0.7558

Now, we will find the firm's beta under new capital structure:

Levered Beta = Unlevered Beta * [1 + {(1 - t) * (D/E)}]

= 0.7558 * [1 + {(1 - 0.40) * (0.5 / 0.5)}]

= 0.7558 * [1 + 0.6] = 0.7558 * 1.6 = 1.21

New cost of equity = Risk-free rate + [Beta * Market Risk Premium]

= 6% + [1.21 * 7%]

= 6% + 8.47%

= 14.47%